# MSRS: Adaptive Multi-Subspace Representation Steering for Attribute Alignment in Large Language Models

### v3

@@ F.2 Subspace Rank $R$
We also explored the effect of subspace rank $R$ on the model’s performance. Based on the rank selection strategy suggested in the ReFT  (Wu et al., 2024a), we tested different values of $R$, including $R=\{4,8,12,16,20\}$, and evaluated the model’s average performance across all datasets on the validation set. In addition to performance metrics, we also reported three key computational efficiency metrics for each $R$ setting: Latency, Throughput, and TFLOPS. These values are also averaged across datasets, ensuring a fair comparison of both performance and efficiency.
@@
1. Performance Trends: As shown in Table 12, $R=8$ consistently outperforms other rank settings in most of the evaluation metrics across multiple datasets. Specifically, MC1 (34.91) and MC2 (56.32) for $R=8$ are the highest among all rank configurations. Although $R=12$ performs slightly better in terms of Sorry-bench score and Helpful scores, it does not show significant improvement in the overall performance compared to $R=8$ . We also observe that performance does not necessarily improve with an increase in $R$. This is likely due to the task-specific needs for representation power. More complex tasks, which require richer and more nuanced representations, benefit from a higher rank, while simpler tasks do not show significant gains with larger ranks. Thus, the optimal choice of $R$ is task-dependent, and $R=8$ strikes the best balance between model capacity and performance for our focused tasks.
2. Computational Cost Consideration: As shown in Table 12, increasing the subspace rank $R$ improves representational capacity but also raises computational costs. Latency grows from 400 ms at $R=4$ to 710 ms at $R=20$, while throughput drops from 160 to 100 requests/sec and TFLOPS increase from 4.15 to 8.20. Given this trade-off, while $R=12$ can yield higher performance on specific metrics, the best balance between performance and computational cost is achieved with $R=8$. The performance gains from increasing the rank beyond 8 do not outweigh the additional computational burden. Therefore, $R=8$ is the optimal choice as it provides strong performance while maintaining a manageable computational cost.
In summary, our analysis shows that the choice of subspace rank $R$ has a notable impact on both model performance and computational efficiency. The best overall trade-off is achieved with $R=8$, which consistently delivers strong performance across datasets while maintaining manageable computational cost.
@@ Appendix G Analysis of Shared Subspace and Singular Value Energy
In our approach, the shared subspace captures common activation directions across attributes, facilitating better integration. We analyze the relationship between the shared subspace ratio and performance by examining the energy captured by the singular vectors in $\Sigma_{c}$. We observe that the optimal shared subspace ratio corresponds to the point where the cumulative energy from the top singular values increases most significantly. This ensures that the shared subspace effectively captures cross-attribute correlations while maintaining sufficient space for attribute-specific subspaces to preserve their respective features. This balance ensures that the attribute-specific subspaces retain sufficient capacity to capture their unique directions, leading to improved overall performance. Based on our analysis, we choose the shared subspace size that accounts for at least 60% of the total energy in $\Sigma_{c}$, optimizing performance across attributes. This strategy outperforms methods that divide the space equally among subspaces, demonstrating the effectiveness of energy-based selection for shared subspaces.

### v4

@@ F.2 Subspace Rank $R$
We also explored the effect of subspace rank $R$ on the model’s performance. Based on the rank selection strategy suggested in the ReFT  (Wu et al., 2024), we tested different values of $R$, including $R=\{4,8,12,16,20\}$, and evaluated the model’s average performance across all datasets on the validation set. In addition to performance metrics, we also reported three key computational efficiency metrics for each $R$ setting: Latency, Throughput, and TFLOPS. These values are also averaged across datasets, ensuring a fair comparison of both performance and efficiency.
@@
1. Performance Trends: As shown in Table 14, $R=8$ consistently outperforms other rank settings across most evaluation metrics. Specifically, MC1 (34.91) and MC2 (56.32) achieve their peaks at $R=8$. While $R=12$ shows marginal gains in Sorry-bench and Helpfulness, it fails to deliver a significant overall improvement. Notably, we observe that performance does not linearly scale with an increase in $R$. In our framework, a larger $R$ is inherently tied to an increased capacity for handling more attributes or larger attribute-specific subspaces. However, expanding $R$ blindly can lead to subspace crowding, where the shared subspace becomes dominated by dominant attributes, causing interference that degrades the precision of attribute-specific steering. This phenomenon aligns with our observations in Section 5.6, where we analyze the trade-offs between attribute quantity and steering fidelity. Thus, $R=8$ strikes the optimal balance between sufficient representation capacity and minimal cross-attribute interference for the tasks evaluated.
2. Computational Cost Consideration: As shown in Table 14, increasing the subspace rank $R$ raises computational costs. Latency grows from 400 ms at $R=4$ to 710 ms at $R=20$, while throughput drops from 160 to 100 requests/sec and TFLOPS increase from 4.15 to 8.20. Given this trade-off, while $R=12$ can yield higher performance on specific metrics, the best balance between performance and computational cost is achieved with $R=8$.
@@ Appendix G Analysis of Shared Subspace and Singular Value Energy
To better understand the impact of the shared subspace energy threshold on performance, we conducted a sensitivity analysis by varying the threshold from 40% to 80%. We evaluated two models, Llama3-8B-Instruct and Qwen2-7B-Instruct, on three different attribute sets: ”truthful & bbq,” ”alpaca & refusal,” and ”helpful & coherence & verbosity.” For each model and rank setting (with $R=8$ and $R=12$), we measured the performance of the model across various energy thresholds and calculated the average performance for each setting.
Table 15: Sensitivity analysis results for varying shared-energy thresholds in multi-attribute steering. Performance is shown for models at ranks $R=8$ and $R=12$ across different attribute sets.
Model Rank ($R$) Energy Threshold ($\theta$) Truthful & BBQ Alpaca & Refusal Helpful & Coherence & Verbosity
Llama3-8B-Instruct 8 40% 48.78 42.61 3.11
50% 51.37 44.23 3.21
60% 51.91 44.45 3.27
70% 51.85 44.38 3.26
80% 50.13 43.38 3.23
Llama3-8B-Instruct 12 40% 49.09 41.71 3.04
50% 51.21 42.8 3.13
60% 51.64 43.2 3.15
70% 51.57 42.1 3.15
80% 50.72 41.35 3.08
Qwen2-7B-Instruct 8 40% 48.31 42.17 3.16
50% 49.78 43.38 3.27
60% 50.73 44.75 3.31
70% 50.66 44.81 3.31
80% 49.17 44.06 3.22
Qwen2-7B-Instruct 12 40% 47.96 40.51 3.08
50% 49.63 42.74 3.25
60% 50.31 43.96 3.36
70% 50.24 43.76 3.37
80% 49.14 41.65 3.21
From Table 15, we observe that very low thresholds (e.g., 40%) allocate too little capacity to the shared subspace, which prevents effective combination of steering capabilities across different attributes, negatively affecting overall performance in multi-attribute steering. This is evident across all models, where performance at 40% is generally lower than at higher thresholds. On the other hand, very high thresholds (e.g., 70–80%) allocate too much capacity to the shared subspace, which correspondingly shrinks the specific subspaces. This reduces the effective steering capacity available for each individual attribute and leads to weaker control in the multi-attribute setting.
Thresholds in the 50–70% range yield consistently strong performance, with 60% sitting comfortably in the middle of this ”high-performance.” This threshold provides a balanced trade-off between retaining shared knowledge and maintaining attribute-specific control, as evidenced by the steady performance across different models and attribute sets. Therefore, the 60% threshold emerges as a robust default that provides reliable performance across a wide range of settings, offering a good balance between shared and specific subspace allocation.
Appendix H Sensitivity Analysis: From Data Coupling to Representation Collapse
A fundamental question in representation engineering is how data-level coupling transmits to representation-level entanglement. To investigate this, we designed a controlled experiment that simulates varying degrees of dataset contamination, observing the point at which the learned subspace $R$ loses its semantic specificity relative to the target attribute.
Construction of Oracle Subspaces. To establish a ground truth, we first curated “pure” datasets for two attributes: Helpfulness (target) and Coherence (distractor). To ensure purity, we employed a stratified sampling strategy: for the pure Helpfulness set $\mathcal{D}_{help}^{*}$, we selected samples with maximal Helpfulness scores while uniformly sampling across the Coherence spectrum to decorrelate the features. A similar process was applied to generate $\mathcal{D}_{coh}^{*}$. We trained independent ReFT probes on these pure sets to obtain the oracle subspaces $R_{help}^{*}$ and $R_{coh}^{*}$.
Contamination Experiment. We then constructed mixed datasets $\mathcal{D}_{mix}(\alpha)$ by blending the pure sets with a mixing ratio $\alpha$, representing the proportion of distractor data:
$\mathcal{D}_{mix}(\alpha)=(1-\alpha)\cdot\mathcal{D}_{help}^{*}+\alpha\cdot\mathcal{D}_{coh}^{*}$ (7)
We trained a steering matrix $R_{mix}$ on each $\mathcal{D}_{mix}(\alpha)$ and evaluated two metrics:
  1. 1.
    Target Performance: The steering efficacy on Helpfulness relative to the baseline.
  2. 2.
    Relative Entanglement Score (RES): Unlike absolute similarity, we propose a relative metric that penalizes alignment with the distractor while rewarding fidelity to the target. It is defined as the ratio of projection strengths:
    $\text{RES}(R_{mix})=\frac{\|R_{mix}(R_{coh}^{*})^{\top}\|_{F}}{\|R_{mix}(R_{help}^{*})^{\top}\|_{F}+\epsilon}$ (8)
    where a lower RES indicates that the learned subspace captures dominant target semantics with minimal distractor leakage. An RES >1>1 implies the subspace is more aligned with the noise than the signal.
Results and Discussion. Figure 8 illustrates the trajectory of performance and entanglement as the contamination ratio $\alpha$ increases from $20\%$ to $80\%$. We observe a distinct non-linear relationship:
  • Robustness Regime ($\alpha\leq 0.55$): The method demonstrates significant resilience. Even when the distractor data constitutes the majority ($\alpha=0.55$), the RES remains relatively low (<0.62<0.62), indicating that the optimization landscape still favors the distinct structural features of the target attribute over the distractor. Consequently, the Target Performance remains stable ($12.5\%$), proving that MSRS can effectively filter out “data noise” as long as the target signal remains discriminative.
  • The Breakdown Point ($\alpha\approx 0.61$): A critical phase transition occurs at $\alpha=0.61$. Here, the RES surges to $0.83$, approaching unity. This suggests a “semantic tipping point” where the distractor gradients begin to dominate the subspace formation, diluting the target direction. This correlates perfectly with the observed sharp drop in steering performance.
  • Residual Efficacy: Notably, even at $\alpha=0.80$ (highly contaminated), the performance gain remains positive ($1.5\%$). This confirms that even under severe coupling, the learned subspace retains a non-trivial projection onto $R_{help}^{*}$, validating the robustness of our subspace-based intervention.
Appendix I Fit Parameters for Scaling Analysis
We provides the parameters for the fits used in the scaling analysis of Figure 10. The fits model the relationships between attribute count $n$ and three metrics: rank $k$, average performance $p$, and interference $i$. The functional forms are:
  • Linear fit for rank: $k(n)=an+b$.
  • Quadratic fit for performance: $p(n)=an^{2}+bn+c$.
  • Cubic fit for interference: $i(n)=an^{3}+bn^{2}+cn+d$.
The parameters and goodness-of-fit ($R^{2}$) values are summarized in Table 16, followed by the explicit equations.
Table 16: Parameters and $R^{2}$ values for fits in Figure 10.
Metric Fit Type $a$ $b$ $c$ $d$ $R^{2}$
Rank ($k$) Linear 2.3091 3.1818 0.9427
Performance ($p$) Quadratic -0.0009 0.0024 0.6735 0.9520
Interference ($i$) Cubic -0.0012 0.0287 -0.0861 0.0800 0.9903
The fitted equations are:
$k(n)$ $=2.3091n+3.1818,$ (9)
$p(n)$ $=-0.0009n^{2}+0.0024n+0.6735,$ (10)
$i(n)$ $=-0.0012n^{3}+0.0287n^{2}-0.0861n+0.0800.$ (11)
These parameters quantify the scaling behavior of multi-attribute steering, with high $R^{2}$ values indicating robust fits, though the limited range ($n\leq 10$).
Appendix J Related Works
Activation Steering Methods. Activation steering aims to adjust activations in specific layers or neurons to guide the model’s output towards desired attributes, without modifying its parameters (Im and Li, 2025). Various approaches have been developed recently (Cao et al., 2024; Bayat et al., 2025; Oozeer et al., 2025). For example, Contrastive Activation Addition (CAA) (Rimsky et al., 2024) computes steering vectors by averaging activation differences between positive and negative examples, which are then added to token positions during inference to control model behavior. Inference-Time Intervention (ITI) (Li et al., 2023a) shifts model activations during inference along predefined directions across attention heads, improving the truthfulness of LLMs. ACT (Wang et al., 2025) trains multiple steering probes on different steering vectors determined by clustering, obtaining steering vectors for different steering patterns. MAT-Steer uses orthogonal constraints to train activation steering vectors, thereby reducing conflicts between steering directions for different attributes (Nguyen et al., 2025). However, previous methods primarily address steering for individual attributes or rely on simple combinations of steering vectors. To solve these limitations, we focus on mitigating interference and optimizing composability across multiple attributes.
Representation Fine-Tuning Methods. In these methods, models will be steered through representation editing. Unlike methods that apply one-rank steering vectors, these approaches extend it by using higher-rank matrices and enhance the expressive power of steering vectors, allowing for richer control over model behavior (Wu et al., 2024). Localized Fine-Tuning (LoFIT) (Yin et al., 2024) identifies critical attention heads for a task and trains offset vectors to modify their hidden representations, offering targeted adjustments. Compositional Subspace Representation Fine-Tuning (CS-ReFT) (Zhou, 2025) advances this by learning orthonormal subspace transformations for distinct skills, composed via a lightweight router, isolating edits in the hidden state to minimize cross-task interference. Unlike previous methods that train steering functions in the same space, we aim to develop representation fine-tuning methods to tune different attribute-specific subspaces and achieve the adaptive integration of multiple attribute steering spaces.